OpenStudy (anonymous):

Find the equation of the line through (-4,0) and (0,2). Give it in slope-intercept form.

OpenStudy (anonymous):

first, find the slope of the line through (-4,0) and (0,2) using the slope formula: \(\large m=\frac{y_2-y_1}{x_2-x_1}= \) ???? what's the slope?

OpenStudy (anonymous):

It doesn't give me the slope

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

use the coordinates of the given points to calculate the slope...

OpenStudy (anonymous):

ok... slope is 1/2. Do you the y-intercept?

OpenStudy (anonymous):

*know

OpenStudy (anonymous):

look carefully at the points given to you...

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

correct...

OpenStudy (anonymous):

a line in slope-intercept form is given as: \(\large y=mx + b \) where m is your slope and b is your y-intercept. all you need to do now is replace the the values you found into that equation.

OpenStudy (anonymous):

So it's y=1/2x+2?

OpenStudy (anonymous):

yes... u got it.... :)

OpenStudy (anonymous):

Thank you! :DD